Section 02  Algebra

Section 02: Algebra

Invited Lectures

Friday, 21 August  MA 005

16:15 – 17:00 William M. Kantor, University of Oregon, Eugene, USA Simple Groups in Computational Group Theory (I: p.41) Chair: Charles C. Sims

17:15 – 18:00 Aleksandr Pukhlikov, Institute for Systems Analysis, Moscow, Russia Birational automorphisms of algebraic varieties (I: p.42) Chair: Victor Vadimovich Batyrev

Saturday, 22 August  MA 005

16:15 – 17:00 Eric Mark Friedlander, Northwestern University, Evanston, USA Geometry of Infinitesimal Group Schemes (I: p.36) Chair: Uwe Jannsen

17:15 – 18:00 Gunter Malle, Universität Heidelberg, Heidelberg, Germany Spetses (I: p.41) Chair: John S. Wilson

Monday, 24 August  MA 005

14:00 – 14:45 Sergei Ivanov, University of Illinois, Urbana, USA On the Burnside problem for groups of even exponent (I: p.38) Chair: Otto H. Kegel

15:00 – 15:45 Idun Reiten, Norwegian University of Science and Technology, Trondheim, Norway Tilting Theory and Quasitilted algebras (I: p.43) Chair: Claus Michael Ringel

Tuesday, 25 August  MA 005

16:15 – 17:00 Aner Shalev, The Hebrew University, Jerusalem, Israel Simple groups, permutation groups, and probability (I: p.45) Chair: Bernhard H. Neumann

17:15 – 18:00 Jeremy C. Rickard, University of Bristol, Bristol, England The abelian defect group conjecture (I: p.45) Chair: Dieter Happel

Algebra  Section 02

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Rings  MA 042

Chair: Toma Albu

14:00 – 14:15 Hyoung J. Ko, Yonsei University, Seoul, Korea, Republic of Some complete intersection and their minimal prime ideals (p.20)

14:15 – 14:30 Takasi Nagahara, Okayama Univ., Okayama-Ken, Japan Classification of rings satisfying some generalized Herstein’s condition (p.24)

14:30 – 14:45 Sylvia Chin-Pi Lu, University of Colorado, Denver, Colorado, USA The Zariski topology on the prime spectrum of a module (p.22)

Algebras I  MA 042

Chair: Alexandre V. Mikhalev

15:00 – 15:15 Askar Dzhumadil’daev, Institute of Mathematics, Almaty, Kazakhstan Cohomologies of right-symmetric algebras and Novikov algebras (p.17)

15:15 – 15:30 Ekaterina A. Vasilieva, Moscow State University, Moscow, Russia Involutive Bases of Free Algebras (p.31)

15:30 – 15:45 Alexander A Totok, Moscow State University, Moscow, Russia Actions of Hopf algebras on noncommutative algebras

Semigroups  MA 311

Chair: Hans-Jürgen Vogel

16:15 – 16:30 Olga Macedonska, Silesian Technical University, Gliwice, Poland On semigroup identities in groups (p.22)

16:30 – 16:45 Boris M. Schein, University of Arkansas, Fayetteville, AR, USA A new construction for free inverse semigroups (p.28)

16:45 – 17:00 Vučić Dašić, Univerzitet Crne Gore, Podgorica, Yugoslavia Hyperrings and hypernear-rings of the multivalied mappings (p.16)

Topological Algebra  MA 042

Chair: Karl-Heinrich Hofmann

16:15 – 16:30 Nouzha El Yacoubi, Dept. Maths & Info. Faculty of Sciences, Rabat, Morocco J-Generalized topological divisors of zero in topological Jordan algebras (p.18)

16:30 – 16:45 Rumara Jewett, SUNY-Albany, Dept. of Mathematics, Albany, NY, USA Almost Convexity of a Certain Nilpotent Group (p.19)

16:45 – 17:00 Mary Nailevna Sabitova, Kazan State University, Kazan, Russia The automorphisms of affine spaces and polynomial manifolds (p.27)

Groups I  MA 311

Chair: Wujie Shi

17:15 – 17:30 Herbert Abels, Fakultät für Mathematik, Universität Bielefeld, Bielefeld, Germany Reductive groups as metric spaces (p.12)

17:30 – 17:45 Oleg I. Tavgen, Belorussian State University, Minsk, Republic of Belarus The free product of cyclic groups in the group PSL_{2}(ℂ) (p.30)

17:45 – 18:00 Matthias Trittler, Universität Ulm, Ulm, Germany Normalteiler symplektischer Gruppen (p.30)

Semirings/groups and Algebras  MA 042

Chair: Boris M. Schein

17:15 – 17:30 Irena I. Artamonova, Moscow State University, Moscow, Russia We consider semirings with 1 ≠ 0 (p.12)

17:30 – 17:45 Godwin Ekhaguere, Association of African Universities, Accra, Ghana Algebraic representation theory of partial algebras (p.17)

17:45 – 18:00 Miroslav, D. Ćirić, University of Niš, Niš, Yugoslavia Atomic decompositions of algebras and structures

Short Communications  Thursday, 20 August
Noetherian Rings  MA 311

Chair: Kurt Rosenbaum

16:15 – 16:30 Vijay Kodiyalam, The Institute of Mathematical Sciences, Chennai, India Asymptotic behaviour of Castelnuovo-Mumford regularity (p.20)

16:30 – 16:45 Syed Tariq Rizvi, The Ohio State University, Lima, USA On noetherian and artinian rings (p.26)

16:45 – 17:00 Sylvia M. Wiegand, University of Nebraska, Lincoln, USA Building Noetherian domains using completions (p.31)

Rings and Modules I  MA 311

Chair: Sylvia M. Wiegand

17:15 – 17:30 Toma Albu, University of Wisconsin-Milwaukee, Milwaukee, Wisconsin, USA Global Krull dimension and global dual Krull dimension of rings (p.12)

17:30 – 17:45 Jilyana Cazaran, University of Tasmania, Hobart, Australia Sufficient conditions for finite commutative rings to be principal ideal rings and an application to error-correcting codes (p.14)

17:45 – 18:00 Reinhard Börger, FernUniversität Hagen, Hagen, Germany Infinitary Linear Combinations in Cotorsion Modules (p.14)

Short Communications  Friday, 21 August
Groups II  MA 311

Chair: Mikhail H. Klin

14:00 – 14:15 Hubert Kiechle, Zentrum Mathematik, TU München, München, Germany Frobenius groups with many involutions (p.20)

14:15 – 14:30 Anatolii S. Kondratiev, Institute of Mathematics and Mechanics, Ekaterinburg, Russia Finite linear groups of small degree (p.20)

14:30 – 14:45 Wujie Shi, Southwest-China Normal University, Chongqing, China, Peoples Republic of A new classification of groups (p.28)

Groups III  MA 311

Chair: Andrew J. Woldar

15:00 – 15:15 Vahagn H. Mikaelian, Math. Institut, Uni. Würzburg, Würzburg, Germany Embeddings of countable groups into 2-generator groups with additional properties (p.31)

15:15 – 15:30 Igor Ya. Subbotin, National University, Inglewood, CA, USA On groups with many periodic factor-groups (p.30)

15:30 – 15:45 Leonid A. Kurdachenko, University of Dnepropetrovsk, Dnepropetrovsk, Ukraine On some infinite dimensional linear groups (p.21)

Short Communications  Saturday, 22 August
Representations of Groups  MA 311

Chair: Alexander S. Kleshchev

14:00 – 14:15 Mohammad Reza Darafsheh, University of Tehran, Tehran, Iran The character table of the group GL_{2}(q) when extended by a certain group of order two (p.15)

14:15 – 14:30 Anna Helversen-Pasotto, Laboratoire “Jean-Alexandre Dieudonné”, Nice, France Représentations de GL(n, F_{q}) et sommes de Gauss (p.18)

14:30 – 14:45 Alexei Davydov, Moscow State University, Department of Math., Moscow, Russia Drinfel’d theory and representations of finite groups (p.16)

Commutative Rings  MA 311

Chair: Peter Draexler

15:00 – 15:15 Alexander Levin, The Catholic Univ. of America, Washington, D.C., USA Multivariable dimension polynomials and their invariants (p.21)

15:15 – 15:30 Venkat Murali, Rhodes University, Grahamstown, South Africa On Krull Intersection Theorem of Fuzzy Ideals (p.24)

15:30 – 15:45 Cornelia G. Naude, University of South Africa, Pretoria, South Africa Asymptotic behaviour of attached primes (p.25)

Short Communications  Monday, 24 August
Algebras II  MA 311

Chair: M. Susan Montgomery

16:15 – 16:30 Mihai I. Cipu, Institute of Mathematics, Bucharest, Romania Multiplicities and pure resolutions for Gorenstein codimension four algebras (p.15)

16:30 – 16:45 Michael Nüsken, Universität Konstanz, Konstanz, Germany Unique Tensor Factorization of Algebras (p.26)

16:45 – 17:00 Andrej A. Zolotykh, Moscow State University, Moscow, Russia Standard bases for ideals of free algebras over commutative rings (p.32)

Hopf Algebras  MA 311

Chair: Alexander A. Mikhalev

17:15 – 17:30 Sorin Dascalescu, University of Bucharest, Bucharest, Romania Pointed Hopf algebras by Ore extensions and classification results (p.16)

17:30 – 17:45 M. Susan Montgomery, University of Southern California, Los Angeles, CA, USA Prime Ideals in Hopf Galois Extensions (p.24)

17:45 – 18:00 Peter Schauenburg, Universität München, München, Germany Hopf algebras with a weak projection (p.27)

Short Communications  Tuesday, 25 August
Rings and Modules II  MA 311

Chair: Christine Bessenrodt

14:00 – 14:15 Sergey P. Strunkov, Dept. of Crptol. and Discr. Math., Moscow, Russia On p-blocks of minimal defect in a finite group and Gelfand pairs (p.29)

14:15 – 14:30 Johannes H. Meyer, University of the Orange Free State, Bloemfontein, South Africa Forcing Linearity Numbers (p.23)

14:30 – 14:45 Hossein Zand, The Open University, Milton Keynes, England Duality in radicals of associative rings (p.32)

Miscellaneous  MA 311

Chair: Corinna Wiedorn

15:00 – 15:15 Martin Goldstern, Technische Universität Wien, Wien, Austria Interpolation of monotone functions in lattices (p.18)

15:15 – 15:30 Ming-Chang Kang, National Taiwan University, Taipei, China, Taiwan Rationality of Some Projective Linear Actions (p.19)

15:30 – 15:45 Vyacheslav A. Artamonov, Moscow State University, Moscow, Russia Quantum polynomial algebra is a multiparameter deformation of an algebra of polynomials, i.e. it is the algebra of regular functions on a quantum affine space (p.12)

Short Communications  Wednesday, 26 August
Algebras III  MA 311

Chair: Shavkat A. Ayupov

15:00 – 15:15 Bernhard Drabant, Department of Applied Mathematics, Cambridge, Great Britain Cross Product Bialgebras (p.17)

15:15 – 15:30 Alexander A. Mikhalev, Moscow State University, Moscow, Russia Free Leibniz algebras (p.23)

15:30 – 15:45 Daouda Sangare, Université de Caen, Hérouville Saint Clair, France Hilbert quasi-polynomial functions and generalized analytic spread (p.27)

Categories  MA 042

Chair: Reinhard Börger

15:00 – 15:15 Hvedri Inassaridze, A. Razmadze Mathematical Institute, Tbilisi, Georgia Karoubi’s Conjecture for polynomial extensions of stable C*-algebras (p.19)

15:15 – 15:30 Fred E.J. Linton, Wesleyan University, New Haven, CT, USA Banach Spaces Monadic over Free Pointed Cones (p.22)

15:30 – 15:45 Siaka Kante Bamba, Université de Cocody, Abidjan, Ivory Coast On quadratic modules (p.13)

Linear Algebra and Quadratic Forms  MA 311

Chair: Stefan Heiss

16:15 – 16:30 Herbert Möller, University of Münster, Münster, Germany New central algorithms in linear algebra (p.23)

16:30 – 16:45 Dilip N. Sheth, Sir Parashurambhau College, Pune, India Completion of a unimodular vector to an invertible matrix (p.28)

16:45 – 17:00 Ulf Rehmann, Universität Bielefeld, Bielefeld, Germany Higher anisotropic kernels of quadratic forms (p.26)

Groups IV  MA 311

Chair: Bettina Wilkens

17:15 – 17:30 Alexei N. Krasilnikov, Moscow Pedagogical State University, Moscow, Russia Identities in metanilpotent groups (p.21)

17:30 – 17:45 Ernest B. Vinberg, Universität Bielefeld, Bielefeld, Germany On a generalization of Coxeter groups

17:45 – 18:00 Munibur Rahman Chowdhury, University of Dhaka, Dhaka, Bangladesh A commutativity theorem for groups (p.15)

Poster Session  Thursday, 20 August 14:00 – 15:45  MA 144

Shavkat A. Ayupov, Institute of Mathematics, Tashkent, Republik of Uzbekistan

Nilpotent Leibniz algebras of low dimension (p.13) Valerii Vaclavovich Benyash-Krivetz, Institute of Mathematics, Minsk, Belarus

On decomposing some finitely generated groups into amalgamated free products (p.13)

Matyas Domokos, Hungarian Academy of Sciences, Budapest, Hungary Determinantal ideals of certain subspace arrangements (p.16)

Kiyomitsu Horiuchi, Konan University, Kobe, Japan An expansion of lattice (p.19)

Andrei Marcus, Babeş-Bolyai, Cluj-Napoca, Romania Derived equivalences and Dade’s conjectures (p.22)

Shigeki Matsumoto, Konan University, Kobe, Japan

Solutions of Sylvester Equations AX−XB = C whose coefficients matrices A and B may have common eigenvalues (p.23)

Marek Niezgoda, Institute of Applied Mathematics, Agric. Univ. of Lublin, Lublin, Poland An unified approach to some majorization inequalities (p.25)

Yasushi Ninomiya, Shinshu University, Matsumoto, Japan

Group algebras and nilpotency indices of their radicals (p.25) Carlos Renteria-Marquez, Instituto Politecnico Nacional, Mexico City, Mexico

The a-invariant of some Reed-Muller codes (p.26) Constatin I. Sonin, Moscow State University, Moscow, Russia

Krull Dimension of Malcev-Neumann Rings (p.29) Christophe Soulé, IHES, Bures-sur-Yvette, France

Filtrations on Higher Algebraic K-theory (p.29) Masamichi Tokizawa, Chuo University, Tokyo, Japan

Set product residue semigroups (p.30) Roman Zapatrin, E&F University, St. Petersburg, Russian Federation

Duality and set representations for partially ordered sets (p.32) Konstantin A. Zubrilin, Moscow State University, Moscow, Russian Federation

The structure theory of algebras that satisfy Capelli identities (p.33)